# Automated Progressive Red Teaming

### v2

@@ Abstract
Ensuring the safety of large language models (LLMs) is paramount, yet identifying potential vulnerabilities is challenging. While manual red teaming is effective, it is time-consuming, costly and lacks scalability. Automated red teaming (ART) offers a more cost-effective alternative, automatically generating adversarial prompts to expose LLM vulnerabilities. However, in current ART efforts, a robust framework is absent, which explicitly frames red teaming as an effectively learnable task. To address this gap, we propose Automated Progressive Red Teaming (APRT) as an effectively learnable framework. APRT leverages three core modules: an Intention Expanding LLM that generates diverse initial attack samples, an Intention Hiding LLM that crafts deceptive prompts, and an Evil Maker to manage prompt diversity and filter ineffective samples. The three modules collectively and progressively explore and exploit LLM vulnerabilities through multi-round interactions. In addition to the framework, we further propose a novel indicator, Attack Effectiveness Rate (AER) to mitigate the limitations of existing evaluation metrics. By measuring the likelihood of eliciting unsafe but seemingly helpful responses, AER aligns closely with human evaluations. Extensive experiments with both automatic and human evaluations, demonstrate the effectiveness of ARPT across both open- and closed-source LLMs. Specifically, APRT effectively elicits 54% unsafe yet useful responses from Meta’s Llama-3-8B-Instruct, 50% from GPT-4o (API access), and 39% from Claude-3.5 (API access), showcasing its robust attack capability and transferability across LLMs (especially from open-source LLMs to closed-source LLMs).
@@
Bojian Jiang1,2, Yi Jing2222Equal contribution., Tianhao Shen1, Tong Wu2, Qing Yang2111Corresponding authors., Deyi Xiong1111Corresponding authors. 1College of Intelligence and Computing, Tianjin University, Tianjin, China 2Du Xiaoman Finance, Beijing, China {jiangbojian, thshen, dyxiong}@tju.edu.cn {jingyi, wutong02, yangqing}@duxiaoman.com
Figure 1: Illustration of APRT. In the training process, the Intention Expanding LLM first generates diverse samples that are relatively easy to jailbreak the Target LLM after intention concealment. For each prompt generated by the Intention Expanding LLM, the Intention Hiding LLM does not transfer its original intention and transforms it to multiple effective samples with deceptive behavior towards the Target LLM. The Target LLM dedicates to generating safe responses to resist the attacks from the Intention Hiding LLM. Two Reward LLMs provide a bias to select new incremental training samples for the Intention Hiding LLM. To swiftly enhance the capability of concealing the intentions within input prompts, the Intention Hiding LLM employs an active learning strategy to prioritize selecting samples whose intention are more challenging to obscure yet successfully elicit responses from the Target LLM that are unsafe yet helpful.
@@ 1 Introduction
Red teaming, a widely used technique for vulnerability discovery in various systems, has become a key approach for assessing and enhancing LLM safety. Manual red teaming relies on human experts to meticulously craft adversarial prompts, aiming to elicit unsafe responses from the Target LLM walkerspider (2022); Bai et al. (2022); Touvron et al. (2023); Wei et al. (2023); Yuan et al. (2024); Kang et al. (2024). While effective, this manual process is expensive, time-consuming, and ultimately unscalable.
@@
In contrast, automated red teaming presents a promising alternative, leveraging computational methods to generate adversarial prompts automatically Perez et al. (2022); Zou et al. (2023); Chao et al. (2024); Liu et al. (2024). This approach has garnered significant interest due to its potential for improved scalability and efficiency. Existing automated red teaming methods fall into two main categories: (1) Template-based methods aiming to develop universal templates which, when combined with raw red teaming instructions, can “jailbreak" target LLMs and elicit unsafe responses Zou et al. (2023); Schulhoff et al. (2023); Shen et al. (2024); Liu et al. (2024); (2) Generation-based approaches that leverage LLMs or LLM-based systems to exploit vulnerabilities in the Target LLM, often employing iterative methods for prompt generation Mehrotra et al. (2023); Chao et al. (2024); Ding et al. (2024); Zeng et al. (2024). While promising, current generation-based red teaming research has not fully harnessed the learning capabilities of parameterized LLMs. Some work Mehrotra et al. (2023); Chao et al. (2024); Ding et al. (2024) constrain the efficacy of LLMs in red teaming framework, using solely fixed-parameter LLMs to launch attacks to the Target LLMs. Our research indicates that progressively adjusting the attack directions based on feedback from the Target LLM is crucial [^1] In Section 5.1, we elucidate the critical role that progressive training plays in automated red teaming through a comprehensive visual analysis.. However, Perez et al. (2022) and Zeng et al. (2024) employ trainable LLMs to attack the Target LLM without using this critical mechanism. The most related to our work is MART [^2] Additional significant differences between APRT and MART are present in Section 3.3 Ge et al. (2023), which adopts a multi-round adversarial training process to improve the efficiency of the Red LLM. The safety of the Target LLM in MART is notably deficient, thereby reducing the difficulty of attacking process. Furthermore, MART’s indiscriminate selection of successful attack samples as incremental training data fails to provide the Red LLM with valuable guidance in attack directions.
@@
To address this critical gap, we propose APRT, Automated Progressive Red Teaming. APRT leverages three integral modules: an Intention Expanding LLM, an Intention Hiding LLM, and an Evil Maker. These three modules work synergistically in an iterative process, progressively attacking the Target LLM and enhancing the ability to conceal malicious intentions within deceptive prompts. As illustrated in Figure 1, APRT leverages active learning techniques to prioritize challenging samples, further boosting its effectiveness. Our experiments on the open-source AdvBench Harmful Behaviors dataset Zou et al. (2023), using both automatic and human evaluations, demonstrate APRT’s superior performance compared to strong baselines. We find that APRT notably induces 54% unsafe yet helpful responses from Llama-3-8B-Instruct. The rates of unsafe yet helpful responses induced from the top-2 closed-source LLMs GPT-4o and Claude-3.5 (both evaluated through API access) are 50% and 39% respectively. These results provide compelling evidence that APRT can effectively transfer attack capabilities to closed-source LLMs by employing samples trained on open-source LLMs.
@@
       $\mathcal{P}^{i}_{\rm{gen}}\leftarrow\mathcal{M}_{\rm{exp}}(\mathcal{P}^{i}_{\rm{att}})$ // expand the attacking prompts
@@ 2 Related Work
Another line of research focuses on meticulously training a LLM or developing a LLM-based system to effectively trigger the Target LLM. PAIR Chao et al. (2024) utilizes a LLM-based attacker to generate improved prompts iteratively. TAP Mehrotra et al. (2023) adopts tree-of-thought technique to generate adversarial prompts. Ge et al. (2023), Ding et al. (2024) and Zeng et al. (2024) meticulously construct an attack framework utilizing LLMs to obscure the explicit intentions of prompts. These sophisticated design aim to effectively circumvent the defensive mechanisms of the Target LLM. The majority of methods in this line has yet to effectively convert red teaming into an effectively learnable task. To address this limitation, we introduce Automated Progressive Red Teaming (APRT), which demonstrates an effective attack capabilities on both aligned open- and closed-source LLMs.
@@ 3 APRT
The entire framework is illustrated in Figure 1. We elaborate this framework in this section. Specifically, we first introduce how to combine the Intention Expanding LLM, the Intention Hiding LLM and the Evil Maker to progressively discover vulnerabilities for the Target LLM and evolve APRT’s attack ability in an automated and iterative manner. Second, we describe the working process of various components in APRT, including Red LLM (including the Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ and Intention Hiding LLM $\mathcal{M}_{\rm{hid}}$), Evil Maker $\mathcal{E}_{\rm{m}}$, Target LLM $\mathcal{M}_{\rm{tgt}}$ and Reward LLMs (safety Reward LLM $\mathcal{R}_{\rm{s}}$ and helpfulness Reward LLM $\mathcal{R}_{\rm{h}}$).
@@ 3.1 Progressive Attack Process
In APRT, the Red LLM identifies vulnerabilities from the Target LLM through automated multi-round attack and progressive evolution process. The process of each iteration is repeated as illustrated in Algorithms 1. The main steps can be displayed as follows:
@@ 3.2 Components in APRT
We initialize Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ with our manually constructed instruction dataset $\mathcal{D}_{\rm{exp}}$ whose input and output prompts have similar words/characters but different semantics. $\mathcal{M}_{\rm{exp}}$ does not update parameters with the multi-round training process. In the attacking process of round $i$, $\mathcal{M}_{\rm{exp}}$ transforms dataset $\mathcal{P}^{i}_{\rm{att}}$ to $\mathcal{P}^{i}_{\rm{gen}}$.
@@
The primary function of Evil Maker $\mathcal{E}_{\rm{m}}$ is to manage the diversity between extended prompts and deceptive prompts by utilizing BLEU scores Papineni et al. (2002) and filter out samples that cannot be successfully rejected by $\mathcal{M}_{\rm{tgt}}$. In the attacking process of round $i$, $\mathcal{E}_{\rm{m}}$ transforms dataset $\mathcal{P}^{i}_{\rm{gen}}$ to $\mathcal{P}^{i}_{\rm{suc}}$. The comprehensive filter process is outlined in Algorithm 4.
@@ 4.1 Data and Models
In order to construct an initialized dataset for the Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ to learn the task of intention expansion, we extracted dataset $\mathcal{D}_{\rm{exp}}$ which consists of 10,000 sentence pairs of the input and output prompts exhibit literal similarity but differ in semantics. Simultaneously, we randomly sampled data from the 14 distinct categories in $\mathcal{P}_{\rm{att}}$ for manual intention obfuscation to jailbreak GPT-4o-mini [^4] https://chatgpt.com/. We retained the pairs (an originally malicious prompt, an intention concealment prompt) whose hidden intention prompts can successfully bypass the defenses of GPT-4o-mini. Ultimately, this process results in a dataset $\mathcal{D}^{0}_{\rm{hid}}$ consist of 300 samples, which we used as seed data to initialize the Intention Hiding LLM $\mathcal{M}^{0}_{\rm{hid}}$.
Evaluation Set
@@
In order to compare the effectiveness of Red Team methods, we employed the vicuna-7b-v1.5 Chiang et al. (2023), Llama-2-7b-chat Touvron et al. (2023) and Llama-3-8B as open-source LLMs for evaluating our methods. In addition, to verify the generality of ARPT, we also used a close-sourced LLMs (GPT-4o-2024-05-13 and claude-3-5-sonnet-20240620 API) to test the performance of APRT.
@@ 4.2 Baseline and Setup
Figure 2: The observed trend in AER (Attack Effectiveness Rate) metric with respect to the number of progressive learning iterations for various open-source Target LLMs
@@ 4.3 Main Results
Table 2: Attack transferability. APRT train different versions of Red LLMs on coresponding Target LLMs. Subsequently, the trained Red LLMs are employed to launch attacks on the Transfer LLMs, followed by the computation of AER metric. * denotes a white-box scenario.
@@
Figure 3: During the training phase of APRT, we systematically visualize the semantic representations generated by both the Intention Expanding LLM and the Intent Hide LLM. This visualization facilitates a comparative analysis of the initial states and epoch-4 checkpoints. In our visualizations, blue dots denote samples where attacks are not successful, whereas red dots indicate samples where attacks are successful.
@@
Figure 5: Comparison of data selection algorithms between MART Ge et al. (2023) and APRT (Ours, active learning-based method)
@@ 6 Conclusion
7 Limitations
We have meticulously designed the ARRT framework, which comprises multiple parameterized and learnable LLMs to establish a robust red team system. This demonstrates the substantial potential of learnable LLMs in red team applications. Nevertheless, the present framework relies on Reward LLMs to guide the progressive training process. As the sophistication of attack prompts increases, the accuracy of Reward LLMs in assessing the safety and helpfulness of Target LLMs’ responses tends to decline. Additionally, developing highly effective Reward LLMs is often cost-prohibitive. To further enhance the capabilities of ARRT, we intend to explore the use of weakly supervised signals to guide the progressive learning process in future work.
8 Ethical Considerations
@@ Appendix A Additional Algorithms, Tables, Figures
Expanded attacking prompt set $\mathcal{P}^{i}_{\rm{exp}}$
@@ 1 $\mathcal{D}^{i}_{\rm{hid}}\leftarrow\{\}$
2 for $\mathcal{P}^{ij}_{\rm{res}}\in\mathcal{P}^{i}_{\rm{res}}$  do
@@
4 for $(s,h,r)\in(\mathcal{P}^{i}_{\rm{suc}},\mathcal{P}^{i}_{\rm{hid}},\mathcal{P}^{ij}_{\rm{res}}$)  do ${s}^{s}\leftarrow\mathcal{R}_{\rm{s}}(s,r)$

### v3

@@ Abstract
Ensuring the safety of large language models (LLMs) is paramount, yet identifying potential vulnerabilities is challenging. While manual red teaming is effective, it is time-consuming, costly and lacks scalability. Automated red teaming offers a more cost-effective alternative, automatically generating adversarial prompts to expose LLM vulnerabilities. However, in current efforts, a robust framework is absent, which explicitly frames red teaming as an effectively learnable task. To address this gap, we propose Automated Progressive Red Teaming (APRT) as an effectively learnable framework. APRT leverages three core modules: an Intention Expanding LLM that generates diverse initial attack samples, an Intention Hiding LLM that crafts deceptive prompts, and an Evil Maker to manage prompt diversity and filter ineffective samples. The three modules collectively and progressively explore and exploit LLM vulnerabilities through multi-round interactions. In addition to the framework, we further propose a novel indicator, Attack Effectiveness Rate (AER) to mitigate the limitations of existing evaluation metrics. By measuring the likelihood of eliciting unsafe but seemingly helpful responses, AER aligns closely with human evaluations. Extensive experiments with both automatic and human evaluations, demonstrate the effectiveness of APRT across both open- and closed-source LLMs. Specifically, APRT effectively elicits 54% unsafe yet useful responses from Meta’s Llama-3-8B-Instruct, 50% from GPT-4o (API access), and 39% from Claude-3.5 (API access), showcasing its robust attack capability and transferability across LLMs (especially from open-source LLMs to closed-source LLMs). The code and seed data are available at https://github.com/tjunlp-lab/APRT.
@@
Bojian Jiang1,2111Equal contribution., Yi Jing2111Equal contribution., Tianhao Shen1, Tong Wu2, Qing Yang2222Corresponding authors., Deyi Xiong1222Corresponding authors. 1College of Intelligence and Computing, Tianjin University, Tianjin, China 2Du Xiaoman Finance, Beijing, China {jiangbojian, thshen, dyxiong}@tju.edu.cn {jingyi, wutong02, yangqing}@duxiaoman.com
Figure 1: Illustration of APRT. In the training process, the Intention Expanding LLM first generates diverse samples that are relatively easy to jailbreak the Target LLM after intention concealment. For each prompt generated by the Intention Expanding LLM, the Intention Hiding LLM transforms it into multiple effective samples with deceptive behavior towards the Target LLM, without changing the original intention of the prompt. The Target LLM dedicates to generating safe responses to resist the attacks from the Intention Hiding LLM. Two Reward LLMs provide a bias to select new incremental training samples for the Intention Hiding LLM. To swiftly enhance the capability of concealing the intentions within input prompts, the Intention Hiding LLM employs an active learning strategy to prioritize selecting samples that can successfully elicit unsafe yet helpful responses from the Target LLM with intentions that are difficult to perceive.
@@ 1 Introduction
Red teaming, a widely used technique for vulnerability discovery in various systems, has become a key approach for assessing and enhancing LLM safety. Manual red teaming relies on human experts to meticulously craft adversarial prompts, aiming to elicit unsafe responses from the Target LLM walkerspider (2022); Bai et al. (2022); Touvron et al. (2023); Wei et al. (2023); Yuan et al. (2024); Kang et al. (2024); Wu et al. (2024). While effective, this manual process is expensive, time-consuming, and ultimately unscalable.
@@
In contrast, automated red teaming (ART) presents a promising alternative, leveraging computational methods to generate adversarial prompts automatically Perez et al. (2022); Zou et al. (2023); Chao et al. (2024); Liu et al. (2024). ART has garnered significant interest due to its potential for improved scalability and efficiency. Existing ART methods fall into two main categories: (1) Template-based methods aiming to develop universal templates which, when combined with raw red teaming instructions, can “jailbreak” target LLMs and elicit unsafe responses Shin et al. (2020); Wallace et al. (2021); Zou et al. (2023); Liu et al. (2024); (2) Generation-based approaches that leverage LLMs or LLM-based systems to exploit vulnerabilities in the Target LLM, often employing iterative methods for prompt generation Mehrotra et al. (2023); Chao et al. (2024); Ding et al. (2024); Zeng et al. (2024).
While promising, current generation-based red teaming research has not fully harnessed the learning capabilities of parameterized LLMs. Some studies Mehrotra et al. (2023); Chao et al. (2024); Ding et al. (2024) constrain the efficacy of LLMs in red teaming framework, using solely fixed-parameter LLMs to launch attacks to the Target LLMs. Our research indicates that progressively adjusting the attack directions based on feedback from the Target LLM is crucial. [^1] In Section 5.1, we elucidate the critical role that progressive training plays in automated red teaming through a comprehensive visual analysis. However, Perez et al. (2022) and Zeng et al. (2024) employ trainable LLMs to attack the Target LLM without using this critical mechanism.
The most closely related to our work is MART [^2] Additional significant differences between APRT and MART are present in Section 3.3. Ge et al. (2023), which adopts a multi-round adversarial training process to improve the efficiency of the Red LLM. The safety of the Target LLM in MART is notably deficient, thereby reducing the difficulty of attacking. Furthermore, MART’s indiscriminate selection of successful attack samples as incremental training data fails to provide the Red LLM with valuable guidance in attack directions.
@@
To address this critical gap, we propose APRT, Automated Progressive Red Teaming. APRT leverages three integral modules: an Intention Expanding LLM, an Intention Hiding LLM, and an Evil Maker. These three modules work synergistically in an iterative process, progressively attacking the Target LLM and enhancing the ability of the Red LLM to conceal malicious intentions within deceptive prompts. As illustrated in Figure 1, APRT leverages active learning techniques to prioritize challenging samples, further boosting its effectiveness.
Our experiments on the open-source AdvBench Harmful Behaviors dataset Zou et al. (2023), using both automatic and human evaluations, demonstrate APRT’s superior performance compared to strong baselines. We find that APRT notably induces 54% unsafe yet helpful responses from Llama-3-8B-Instruct. The rates of unsafe yet helpful responses induced from the top-2 (as of our attack assessment) closed-source LLMs GPT-4o and Claude-3.5 (both evaluated through API access) are 50% and 39% respectively. These results provide compelling evidence that APRT can effectively transfer attack capabilities to closed-source LLMs by employing samples trained on open-source LLMs.
@@
       $\mathcal{P}^{i}_{\rm{gen}}\leftarrow\mathcal{M}_{\rm{exp}}(\mathcal{P}_{\rm{att}})$ // expand the attacking prompts
@@ 2 Related Work
Another line of research focuses on meticulously training an LLM or developing an LLM-based system to effectively trigger the Target LLM. PAIR Chao et al. (2024) utilizes an LLM-based attacker to generate improved prompts iteratively. TAP Mehrotra et al. (2023) adopts tree-of-thought technique to generate adversarial prompts. Ge et al. (2023), Ding et al. (2024) and Zeng et al. (2024) meticulously construct an attack framework utilizing LLMs to obscure the explicit intentions of prompts. These sophisticated design aim to effectively circumvent the defensive mechanisms of the Target LLM. The majority of methods in this line has yet to effectively convert red teaming into an effectively learnable task. To address this limitation, we introduce Automated Progressive Red Teaming (APRT), which demonstrates an effective attack capabilities on both aligned open- and closed-source LLMs.
@@ 3 APRT
The entire framework is illustrated in Figure 1. We elaborate this framework in this section. Specifically, we first introduce how to combine the Intention Expanding LLM, the Intention Hiding LLM and the Evil Maker to progressively discover vulnerabilities for the Target LLM and evolve APRT’s attack ability in an automated and iterative manner. Second, we describe the working process of key components in APRT, namely the Red LLM (including both the Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ and Intention Hiding LLM $\mathcal{M}_{\rm{hid}}$), Evil Maker $\mathcal{E}_{\rm{m}}$, Target LLM $\mathcal{M}_{\rm{tgt}}$ and Reward LLMs (safety Reward LLM $\mathcal{R}_{\rm{s}}$ and helpfulness Reward LLM $\mathcal{R}_{\rm{h}}$).
@@ 3.1 Progressive Attack Process
In APRT, the Red LLM identifies vulnerabilities from the Target LLM through automated multi-round attack and progressive evolution process. The process of each iteration is repeated as illustrated in Algorithm 1. The main steps can be displayed as follows:
@@ 3.2 Components in APRT
We initialize Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ with our manually constructed instruction dataset $\mathcal{D}_{\rm{exp}}$ whose input and output prompts have similar words/characters but different semantics. $\mathcal{M}_{\rm{exp}}$ does not update parameters with the multi-round training process. In the attacking process of round $i$, $\mathcal{M}_{\rm{exp}}$ transforms dataset $\mathcal{P}_{\rm{att}}$ to $\mathcal{P}^{i}_{\rm{gen}}$.
@@
The primary function of Evil Maker $\mathcal{E}_{\rm{m}}$ is to manage the diversity between original prompts and extended prompts by utilizing BLEU scores Papineni et al. (2002) and filter out samples that cannot be successfully rejected by $\mathcal{M}_{\rm{tgt}}$. In the attacking process of round $i$, $\mathcal{E}_{\rm{m}}$ transforms dataset $\mathcal{P}^{i}_{\rm{gen}}$ to $\mathcal{P}^{i}_{\rm{suc}}$. The comprehensive filter process is outlined in Algorithm 4.
@@ 4.1 Data and Models
In order to construct an initialized dataset for the Intention Expanding LLM $\mathcal{M}_{\rm{exp}}$ to learn the task of intention expansion, we extracted dataset $\mathcal{D}_{\rm{exp}}$ which consists of 10,000 sentence pairs of the input and output prompts exhibit literal similarity but differ in semantics. Simultaneously, we randomly sampled data from the 14 distinct categories in $\mathcal{P}_{\rm{att}}$ for manual intention obfuscation to jailbreak GPT-4o-mini. [^4] https://chatgpt.com/ We retained the pairs (an originally malicious prompt, an intention concealment prompt) whose hidden intention prompts can successfully bypass the defenses of GPT-4o-mini. Ultimately, this process results in a dataset $\mathcal{D}^{0}_{\rm{hid}}$ consist of 300 samples, which we used as seed data to initialize the Intention Hiding LLM $\mathcal{M}^{0}_{\rm{hid}}$.
Evaluation Dataset
@@
In order to compare the effectiveness of Red Teaming methods, we employed the vicuna-7b-v1.5 Chiang et al. (2023), Llama-2-7b-chat Touvron et al. (2023) and Llama-3-8B as open-source LLMs for evaluating our methods. In addition, to verify the generality of APRT, we also used a close-sourced LLMs (GPT-4o-2024-05-13 and claude-3-5-sonnet-20240620 API) to test the performance of APRT.
@@ 4.2 Baseline and Setup
Figure 2: The observed trend in AER (Attack Effectiveness Rate) metric with respect to the number of progressive learning iterations for various open-source Target LLMs.
@@ 4.3 Main Results
Table 2: Attack transferability. APRT trains different versions of Red LLMs against coresponding Target LLMs. Subsequently, the trained Red LLMs are employed to launch attacks on the Transfer LLMs, followed by the computation of AER metric. * denotes a white-box scenario.
@@
Figure 3: During the training of APRT, we systematically visualize the semantic representations generated by both the Intention Expanding LLM and the Intention Hiding LLM. This visualization facilitates a comparative analysis of the initial states and epoch-4 checkpoints. In our visualizations, blue dots denote samples where attacks are not successful, whereas red dots indicate samples where attacks are successful.
@@
Figure 5: Comparison of data selection algorithms between MART Ge et al. (2023) and APRT (Ours, active learning-based method).
@@ 6 Conclusion
Limitations
We have meticulously designed the APRT framework, which comprises multiple parameterized and learnable LLMs to establish a robust red team system. This demonstrates the substantial potential of learnable LLMs in red team applications. Nevertheless, the present framework relies on Reward LLMs to guide the progressive training process. As the sophistication of attack prompts increases, the accuracy of Reward LLMs in assessing the safety and helpfulness of Target LLMs’ responses tends to decline. Additionally, developing highly effective Reward LLMs is often cost-prohibitive. To further enhance the capabilities of APRT, we intend to explore the use of weakly supervised signals to guide the progressive learning process in future work.
Ethical Considerations
@@
Acknowledgments
The present research was partially supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). We would like to thank the anonymous reviewers for their insightful comments.
@@ Appendix A Additional Algorithms, Tables, Figures
Original attacking prompt set $\mathcal{P}_{\rm{att}}$
Expanded attacking prompt set $\mathcal{P}^{i}_{\rm{gen}}$
@@ 1 $\mathcal{D}^{i}_{\rm{hid}}\leftarrow\{\}$
2 for $(\mathcal{P}^{ij}_{\rm{hid}},\mathcal{P}^{ij}_{\rm{res}})\in(\mathcal{P}^{i}_{\rm{hid}},\mathcal{P}^{i}_{\rm{res}})$  do
@@
4 for $(s,h,r)\in(\mathcal{P}^{i}_{\rm{suc}},\mathcal{P}^{ij}_{\rm{hid}},\mathcal{P}^{ij}_{\rm{res}}$)  do ${s}^{s}\leftarrow\mathcal{R}_{\rm{s}}(s,r)$